# Integrability of open boundary driven quantum circuits

### v2

@@ 1 Introduction
Integrability in boundary driven diffusive systems
@@
The initial motivation for this paper was to explore any potential connection between the matrix product form of the NESS and the Yang-Baxter integrability structures of the entire spin chain (boundary+bulk). This motivates us to study the boundary reflection algebra for an $R$ matrix of the factorized form $R(u)=r(u)\otimes r(-u)$ (see below).
@@ 2.2 Two-step Floquet dynamics
For completeness, we first review the construction of the integrable circuit for the cases of periodic and open boundary conditions, in section 2.2.1 and 2.2.2.
@@ 3.1 Building the quantum circuit
For the reason that will be clarified in the following, we construct a quantum circuits where each gate is composed by the tensor product of two copies of the $r(u)$ matrix, $R(u)=r(u)\otimes r(-u)$.
@@ 5.2.1 Analytical expression of the boundary terms
For the class of models we are considering, except model of section 4.2.2 with $h=1$, the first and second derivative of the transfer matrix do not generate the dynamics. In fact, the first derivative vanishes and the second derivative is proportional to the identity operator. The first non-trivial charge is the derivative of order 3 of the transfer matrix. For completeness, since we were not able to find this expressions in the literature, we give the analytical expression of the first two charges in the appendix C. We derived these expressions, as well as the one for the third derivatives, using Mathematica software and the non-commutative product. The precise expression for the third derivative in a general model is quite lengthy, so it is not included in this paper. However, upon request, we are happy to share the Mathematica notebook with anyone interested. For the model we are considering, many terms are zero and the final expression is very compact.
@@ 5.2.3 XX spin chain with magnetic field
It is worth to mention that for this model, an interesting phenomenon happens. The first non-trivial conserved charge is generated by taking the first derivative of the transfer matrix, however, the analytical expression does not coincide with (26) given by Sklyanin in [45]. The reason is that, in Sklyanin’s derivation, he assumes that taking the limit of the product of matrices as the spectral parameter approaches zero is equivalent to taking the limit of each term in the product individually. However, there can be cases, such as the one we are considering, where this assumption does not hold. In our case, one factor diverges in the limit. Taking the product first regularizes the result. We comment further on this statement in Appendix C.
@@ 6 Conclusion
This result is somehow reminiscent to a similar phenomenon in quantum field theory (QFT) with extended defect lines or with an impurity, [48, 49]. There it was observed that only integrable QFTs with diagonal $S$-matrix in the bulk and factorizable scattering exhibiting both transmission and reflection are those associated to generalized-free theories.
Many interesting open questions need to be addressed. One may introduce the coupling between the two free fermionic (XX) chains either in the form of a Hubbard interaction in the unitary context or dephasing in the open systems framework, and again ask the question about factorizability or Lindbladian form of the boundary terms. Furthermore, we observed that the circuit of Fig. 10 where the gate is the XX model, corresponds to a finite free fermionic model with open boundaries and an interacting impurity in the middle. For a particular choice of the parameters of the impurity, one has $U(1)$ symmetry which allows to study impurity related transport properties of the model. Another interesting direction is to study factorizability of $K$-matrices for higher local Hilbert space dimension.
@@ Appendix A Factorized solutions: motivation
Appendix B Allowed transformations
@@
$u\to h(u),\,\,\text{such\, that}\,\,h(u-v)=h(u)-h(v)$ (B.6)
@@
Appendix C Explicit calculation of the first two derivatives of $t(u)$
@@
$t(u)=\text{tr}_{0}K_{0}^{L}R_{0L}\dots R_{02}R_{01}K_{0}^{R}R_{01}R_{02}\dots R_{0L},$ (C.1)
@@
$\lim_{u\to 0}A_{1}(u)A_{2}(u)\dots A_{2L+2}(u)=\lim_{u\to 0}A_{1}(u)\lim_{u\to 0}A_{2}(u)\dots\lim_{u\to 0}A_{2L+2}(u),$ (C.2)
where $A_{i}(u)$ are the matrices entering in the transfer matrix definition. We remark that this is also the assumption used by Sklyanin in [45] and that, for the case of the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the assumption does not hold. We motivate it in section C.1.
@@
C.1 First derivative $t^{\prime}(0)$
@@
$t^{\prime}(0)=2\text{tr}_{0}K^{L}_{0}(0)H_{Bulk}+\text{tr}_{0}K^{L}_{0}(0){K^{R}_{1}}^{\prime}(0)+2\text{tr}_{0}K_{0}^{L}(0)h_{L0}+\big{(}\text{tr}_{0}{K^{L}_{0}}^{\prime}(0)\big{)}\mathbb{I},$ (C.4)
@@
$H_{Bulk}=\sum_{i=1}^{L-1}h_{i,i+1},$ $H^{R}_{1}=\frac{1}{2}{K_{1}^{R}}^{\prime}(0),$ $H^{L}_{L}=\frac{\text{tr}_{0}K_{0}^{L}(0)h_{L0}}{\text{tr}_{0}K_{0}^{L}(0)}.$ (C.7)
@@
For the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the first derivative of the transfer matrix generates the dynamics. For that model, an interesting phenomenon happens and the expression (C.4) does not hold. In fact, the term $2\text{tr}_{0}K_{0}^{L}(0)h_{L0}$ that comes from taking the derivative of $R_{0L}(u)$ ($\text{lim}_{u\to 0}\text{tr}_{0}K_{0}^{L}(u)R_{0L}^{\prime}(u)R_{0,L-1}(u)\dots$) is divergent. However, adding it to the terms coming from $\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots$ cancels the divergent terms. Since (C.3) holds, we may be tempted to conclude that for the term added, the following identity holds
$\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots=\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u).$ (C.8)
@@
C.2 Second derivative $t^{\prime\prime}(0)$
@@
$2\,\text{tr}_{0}K_{0}^{L}(0)\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{L-2}\{h_{i,i+1},h_{j,j+1}\}.$ (C.9)
We did not find this explicit computation in the literature. In [53], the authors gave the expression of $t^{\prime\prime}(u)$ for the case with $\text{tr}_{0}K_{0}^{L}(0)h_{L,0}\propto\mathbb{I}$ and $\text{tr}_{0}K_{0}^{L}(0)=0$. We checked that, under these assumptions, our expressions reduce to theirs. In that case, the expression of the second derivative, takes the expression (C.6). However, in the absence of these assumptions, the charge cannot be written in a form similar to (C.6). The terms acting in the boundary will have interaction range $2$ and the ones in the bulk range $3$.

### v3

@@ 1 Introduction
Integrability in boundary driven diffusive Lindbladian systems
@@
Motivation of the paper
The initial motivation for this paper was to explore any potential connection between the matrix product form of the NESS and the Yang-Baxter integrability structures of the entire spin chain (boundary+bulk). More specifically, a Lindblad system is considered Yang-Baxter integrable if the generator of the dynamics $\mathcal{L}$ commutes with an infinite number of conserved superoperators generated by a transfer matrix, [17, 18, 19, 20]. In particular, the bulk dynamics governed by the Hamiltonian is related to an $R$-matrix that solves the Yang-Baxter equation, while the boundary terms correspond to $K$-matrices that solve the boundary Yang-Baxter equations. We consider the case where the Hamiltonian is the sum of range 2 operators, $H=\sum_{i=1}^{L}h_{i,i+1}$. As we clarify in sec. 5, it is convenient to express the evolution of the density matrix in a doubled Hilbert space. The bulk evolution corresponds to $i(1\otimes H^{T}-H\otimes 1)$. The $R$-matrix associated to it is $R^{\prime}(u)|_{u=0}\sim P(1\otimes h^{T}-h\otimes 1)$. This motivates us to consider an elementary gate with factorized form $R(u)=r(u)\otimes r(-u)$, where $r(u)$ is the $R$-matrix of a spin 1/2 chain, related to the Hamiltonian density as $r^{\prime}(u)|_{u=0}=p\,h$. We study the boundary reflection algebra $K$ associated with this $R$ matrix, focusing on whether all solutions take a factorized form or not. The first case, where the solution is factorizable, is less interesting as it corresponds to two uncoupled quantum circuits (or spin chains in the continuum time limit), while the second case, which is non-factorizable, is non-trivial. For the non-factorizable solution, we analyze the conditions under which it can be expressed as the dissipator term of a Lindbladian (1), and we compare these results with the existing literature on NESS.
@@ 2.1.2 Open Boundary Condition
In Appendix C, we write down the transformation that can be performed on the $K^{R/L}$ matrices such that they remain solutions of (10)/(11).
@@ 2.2 Two-step Floquet dynamics
For completeness, we first review the construction of the integrable circuit for the cases of periodic and open boundary conditions, in section 2.2.1 and 2.2.2. For completeness, in Appendix B, we also consider twisted boundary conditions.
@@ 3.1 Building the quantum circuit
For the reason that we briefly mentioned in the introduction and that will be clarified in the following, we construct a quantum circuits where each gate is composed by the tensor product of two copies of the $r(u)$ matrix, $R(u)=r(u)\otimes r(-u)$.
@@ 4.2.1 The XX spin chain
We note that since all the models obtained are Yang-Baxter integrable, they are explicitly solvable using the Bethe ansatz or some other techniques from the theory of integrable models, for example the Baxter $Q$-operators. For the case of the algebraic Bethe ansatz applied to quantum circuits with periodic boundary conditions, we refer to [45]. The algebraic Bethe ansatz for open boundary conditions is analyzed in [3] for the case of a homogeneous transfer matrix and diagonal boundaries. The inhomogeneous case is straightforward. The case with non-diagonal boundaries is more complex; however, numerous techniques have been developed to address it (see, for example, [46]). Furthermore, for the particular choice of parameters that corresponds to having a $U(1)$ symmetry, it is also possible to perform the coordinate Bethe ansatz, in a manner similar to [47].
@@ 5.2.1 Analytical expression of the boundary terms
For the class of models we are considering, except model of section 4.2.2 with $h=1$, the first and second derivative of the transfer matrix do not generate the dynamics. In fact, the first derivative vanishes and the second derivative is proportional to the identity operator. The first non-trivial charge is the derivative of order 3 of the transfer matrix. For completeness, since we were not able to find this expressions in the literature, we give the analytical expression of the first two charges in the appendix D. We derived these expressions, as well as the one for the third derivatives, using Mathematica software and the non-commutative product. The precise expression for the third derivative in a general model is quite lengthy, so it is not included in this paper. However, upon request, we are happy to share the Mathematica notebook with anyone interested. For the model we are considering, many terms are zero and the final expression is very compact.
@@ 5.2.3 XX spin chain with magnetic field
It is worth to mention that for this model, an interesting phenomenon happens. The first non-trivial conserved charge is generated by taking the first derivative of the transfer matrix, however, the analytical expression does not coincide with (26) given by Sklyanin in [49]. The reason is that, in Sklyanin’s derivation, he assumes that taking the limit of the product of matrices as the spectral parameter approaches zero is equivalent to taking the limit of each term in the product individually. However, there can be cases, such as the one we are considering, where this assumption does not hold. In our case, one factor diverges in the limit. Taking the product first regularizes the result. We comment further on this statement in Appendix D.
@@ 6 Conclusion
This result is somehow reminiscent to a similar phenomenon in integrable quantum field theory (IQFT) with extended defect lines or with an impurity [^20] We thank Z. Bajnok for pointing out the similarity between the two results., [52, 53]. In IQFT, there exist two types of integrable defects, [52]: topological (purely trasmissive) and non-topological (transmissive and reflective) defects. By using the so-called “folding trick”, the defects can be related to the boundary. In [53], it was shown that the only theories that allows non-topological integrable defects are the free theories (free fermion or free boson). Theories characterized by non-topological defects are related to the factorizability of the boundary S-matrices.
Many interesting open questions need to be addressed. One may introduce the coupling between the two free fermionic (XX) chains either in the form of a Hubbard interaction in the unitary context or dephasing in the open systems framework, and again ask the question about factorizability or Lindbladian form of the boundary terms. Furthermore, we observed that the circuit of Fig. 10 where the gate is the XX model, corresponds to a finite free fermionic model with open boundaries and an interacting impurity in the middle. For a particular choice of the parameters of the impurity, one has $U(1)$ symmetry which allows to study impurity related transport properties of the model. Furthermore, since all the models constructed are Yang-Baxter integrable, it is also possible to obtain their spectrum via one of the Bethe ansatz techniques. Another interesting direction is to study factorizability of $K$-matrices for higher local Hilbert space dimension.
@@ Appendix A Factorized solutions: motivation
Appendix B Twisted boundary conditions
We briefly consider here an intermediate case between open and periodic boundary condition, the twisted or quasi-periodic boundary condition. The transfer matrix takes the expression, [20],[38]
$t(u,\mathbf{w})=\text{tr}_{0}G_{0}R_{0L}(u-w_{L})\dots R_{02}(u-w_{2})R_{01}(u-w_{1}),$ (B.1)
where the twist matrix $G$ satisfies
$[R_{i,j}(u),G\otimes G]=0.$ (B.2)
In the circuit language, one can fix the inhomogeneities as
$w_{1}=w_{3}=w_{5}=\dots=w_{L-1}=-\kappa,$ $w_{2}=w_{4}=\dots=w_{L}=\kappa$ (B.3)
and hence
$t(u)=\text{tr}_{0}G_{0}R_{0L}(u-\kappa)\dots R_{02}(u-\kappa)R_{01}(u+\kappa).$ (B.4)
Evaluating the transfer matrix at $u=\pm\kappa$,
$t(\kappa)=P_{1,L}P_{1,L-1}\dots P_{1,3}P_{1,2}U_{2,3}\dots U_{L-2,L-1}U_{L,1}G_{L},$ (B.5)
$t(-\kappa)=P_{1,L}P_{1,L-1}\dots P_{1,3}P_{1,2}G_{L}U_{L-1,L}^{-1}\dots U_{3,4}^{-1}U_{1,2}^{-1}.$ (B.6)
We can define the dynamic as
$M=t(-\kappa)^{-1}t(\kappa)=U_{1,2}U_{3,4}\dots U_{L-1,L}U_{2,3}U_{4,5}\dots U_{L-2,L-1}G_{L}U_{L,1}G_{L}^{-1},$ (B.7)
and recognize the two steps in the evolution
$\mathbb{U}_{1}=\left(\prod_{k=1}^{\frac{L}{2}}U_{2k-1,2k}\right),\,$ $\mathbb{U}_{2}=\left(\prod_{k=1}^{\frac{L}{2}-1}U_{2k,2k+1}\right)G_{L}^{-1}U_{L,1}G_{L}$ (B.8)
which corresponds to the circuit of Figure 11.
This type of boundary conditions are not independent from the open boundary conditions, but they can be related to it, as explained in [3].
Interestingly, from the point of view of quantum circuits, a dynamic similar to Figure 11 can be obtained from the case of open boundary condition where the length of the spin chain is even, [14]. It is easy to see that for that case
$t(u)=\text{tr}_{0}\Big{(}{K}_{0}^{L}(u)R_{0L}(u\!-\!\kappa)\dots R_{02}(u\!-\!\kappa)R_{01}(u\!+\!\kappa)K_{0}^{R}(u)R_{10}(u\!-\!\kappa)R_{20}(u\!+\!\kappa)\dots R_{L0}(u\!+\!\kappa)\Big{)},$ (B.9)
$t(\kappa)=U_{12}U_{34}\dots U_{L-1,L}{K}_{1}^{L}(\kappa)U_{23}U_{45}\dots U_{L-1,L-2}\text{tr}_{0}\Big{(}K_{0}^{R}(\kappa)U_{0L}\Big{)},$ (B.10)
which is equivalent to the dynamics of Figure 11 with the gate $\tilde{U}_{i,j}={K}_{i}^{L}(\kappa)\text{tr}_{0}\Big{(}K_{0}^{R}(\kappa)U_{0j}\Big{)}$.
Figure 11: Quantum circuit from twisted boundary condition. The gate $\tilde{U}$ is related to $U$ as $\tilde{U}_{i,j}=G_{i}^{-1}U_{i,j}G_{i}$.
Appendix C Allowed transformations
@@
$u\to h(u),\,\,\text{such\, that}\,\,h(u-v)=h(u)-h(v)$ (C.6)
@@
Appendix D Explicit calculation of the first two derivatives of $t(u)$
@@
$t(u)=\text{tr}_{0}K_{0}^{L}R_{0L}\dots R_{02}R_{01}K_{0}^{R}R_{01}R_{02}\dots R_{0L},$ (D.1)
@@
$\lim_{u\to 0}A_{1}(u)A_{2}(u)\dots A_{2L+2}(u)=\lim_{u\to 0}A_{1}(u)\lim_{u\to 0}A_{2}(u)\dots\lim_{u\to 0}A_{2L+2}(u),$ (D.2)
where $A_{i}(u)$ are the matrices entering in the transfer matrix definition. We remark that this is also the assumption used by Sklyanin in [49] and that, for the case of the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the assumption does not hold. We motivate it in section D.1.
@@
D.1 First derivative $t^{\prime}(0)$
@@
$t^{\prime}(0)=2\text{tr}_{0}K^{L}_{0}(0)H_{Bulk}+\text{tr}_{0}K^{L}_{0}(0){K^{R}_{1}}^{\prime}(0)+2\text{tr}_{0}K_{0}^{L}(0)h_{L0}+\big{(}\text{tr}_{0}{K^{L}_{0}}^{\prime}(0)\big{)}\mathbb{I},$ (D.4)
@@
$H_{Bulk}=\sum_{i=1}^{L-1}h_{i,i+1},$ $H^{R}_{1}=\frac{1}{2}{K_{1}^{R}}^{\prime}(0),$ $H^{L}_{L}=\frac{\text{tr}_{0}K_{0}^{L}(0)h_{L0}}{\text{tr}_{0}K_{0}^{L}(0)}.$ (D.7)
@@
For the XX spin chain with magnetic field $h=1$ that we treat in section 5.2.3, the first derivative of the transfer matrix generates the dynamics. For that model, an interesting phenomenon happens and the expression (D.4) does not hold. In fact, the term $2\text{tr}_{0}K_{0}^{L}(0)h_{L0}$ that comes from taking the derivative of $R_{0L}(u)$ ($\text{lim}_{u\to 0}\text{tr}_{0}K_{0}^{L}(u)R_{0L}^{\prime}(u)R_{0,L-1}(u)\dots$) is divergent. However, adding it to the terms coming from $\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots$ cancels the divergent terms. Since (D.3) holds, we may be tempted to conclude that for the term added, the following identity holds
$\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u)R_{0L}(u)R_{0,L-1}(u)\dots=\text{lim}_{u\to 0}\text{tr}_{0}{K_{0}^{L}}^{\prime}(u).$ (D.8)
@@
D.2 Second derivative $t^{\prime\prime}(0)$
@@
$2\,\text{tr}_{0}K_{0}^{L}(0)\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{L-2}\{h_{i,i+1},h_{j,j+1}\}.$ (D.9)
We did not find this explicit computation in the literature. In [57], the authors gave the expression of $t^{\prime\prime}(u)$ for the case with $\text{tr}_{0}K_{0}^{L}(0)h_{L,0}\propto\mathbb{I}$ and $\text{tr}_{0}K_{0}^{L}(0)=0$. We checked that, under these assumptions, our expressions reduce to theirs. In that case, the expression of the second derivative, takes the expression (D.6). However, in the absence of these assumptions, the charge cannot be written in a form similar to (D.6). The terms acting in the boundary will have interaction range $2$ and the ones in the bulk range $3$.